OpenStudy (anonymous):

solve using the quadratic formula: (x-3)(x+1)=-6

OpenStudy (mathstudent55):

Mulitiply out the left side and collect like terms. Then add 6 to both sides to get the -6 on the left side. Then use the quadratic formula.

OpenStudy (anonymous):

thats where i am stuck. i have x^2-2x+9=0

OpenStudy (mathstudent55):

Do foil again on left (x - 3)(x + 1) = -6 x^2 + x - 3x - 3 = -6 Add 6 to both sides: x^2 - 2x + 3 = 0

OpenStudy (anonymous):

ive got that.. then what do i do?

OpenStudy (mathstudent55):

Now use the quadratic formula, with a = 1, b = -2 and c = 3. Do you know the quadratic formula?

OpenStudy (anonymous):

oh okay got it yes thank you!!

OpenStudy (mathstudent55):

\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (mathstudent55):

ok, wlcm